Can the Intermediate Value Theorem be applied to show that there is a root of the
equation X^5 - X³ + 3x - 5 =0 in the given interval [2,1] ? If yes, apply it
1
Expert's answer
2020-03-03T09:15:02-0500

Define the function "f(x)=x^5-x^3+3x-5." The function "f(x)" is continuous on the closed interval "[1, 2]" as polynomial.


"f(1)=1^5-1^3+3(1)-5=-2<0"

"f(2)=2^5-2^3+3(2)-5=25>0"

By the Intermediate Value Theorem "f" must have a zero between "1" and "2."

Hence the Intermediate Value Theorem can be applied to show that there is a root of the

equation "x^5-x^3+3x-5=0" in the given interval "[1,2]."